OpenStudy (anonymous):

6 5/8 - 1/6 = i need help

OpenStudy (anonymous):

Is it \(\frac{65}{8}\) or \(6 \frac{5}{8}\) ??

OpenStudy (anonymous):

I am assuming it to be second one: \(6\frac{5}{8}\) = \(\large \frac{53}{8}\) \[\frac{53}{8} - \frac{1}{6} = \frac{159 - 4}{24} = \frac{155}{24}\]

OpenStudy (anonymous):

not the 65 one the other one idk how to make do the simble thing lol

OpenStudy (anonymous):

i gotta find the defference

OpenStudy (anonymous):

I have found the difference..

ganeshie8 (ganeshie8):

\(\huge 6\frac{5}{8} - \frac{1}{6}\) \(\huge 6 + \frac{5}{8} - \frac{1}{6}\) \(\huge 6 + \frac{5*3}{8*3} - \frac{1*4}{6*4}\) \(\huge 6 + \frac{15}{24} - \frac{4}{24}\) \(\huge 6 + \frac{15-4}{24}\) \(\huge 6 + \frac{11}{24}\) \(\huge 6 \frac{11}{24}\)